OpenStudy (moongazer):

If a 75-kg runner can accelerate from standing start to a speed of 10m/s in 3 seconds, ho much power is generated?

OpenStudy (moongazer):

show your solutions please

OpenStudy (king):

wait

OpenStudy (king):

u can first find acceleration through v=u+at ,u will get time as 10/3,then find displacement by x=(v+u)/2.t u will get displacement as 50/3 m,then work done is W=m.a.d=75.10/3.50/3 that is 12500/3.Then power is W/t that is P=12500/3 divided by 10/3 that is equal to 1250 watts or 1.25 kW

OpenStudy (king):

if u dont understand anything let me know and i will xplain that part again..

OpenStudy (king):

if u understand please give me a message that u hav understood

OpenStudy (moongazer):

did you know that power=work/time

OpenStudy (king):

yeah its a formula

OpenStudy (moongazer):

why did my teacher gave me an answer of 2450 W???

OpenStudy (king):

wait let me check my answer

OpenStudy (king):

correction displacement is 15 m,so work =75.10/3.15 that is 3750 J so power = 3750/10/3 that is 3750 divided by 3 that is equal to 1250 I am getting only 1 answer so i think u should write this down and then ask ure teacher as she might hav been wrong or else if i am wrong please let me know

OpenStudy (moongazer):

why is the displacement 15?

OpenStudy (moongazer):

my teacher's solution is [(735)(10)]/3

OpenStudy (anonymous):

1250

OpenStudy (moongazer):

why?